OpenStudy (anonymous):

2/5 = p/15 what does the p =

OpenStudy (anonymous):

(2/5)*15=p p=6

OpenStudy (anonymous):

You have to multiply first.... 5 * p = 5p 2 * 15 = 30 30/5 can you divided it?

OpenStudy (anonymous):

k thx so much :)

hero (hero):

Cross Multiply: \[\frac{2}{5} = \frac{p}{15}\] 5p = 30 p = 30/5 p = 6